OpenStudy (anonymous):

Solve the following equation by clearing fractions -2 +3/4x = 5x-1/3

OpenStudy (anonymous):

So you want to get rid of the fractions, multiply everything by the LCD, which would be 12, and work from there.

OpenStudy (mertsj):

\[-2+\frac{3}{4x}=5x-\frac{1}{3}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

That's the problem but the x is behind the whole fraction 3/4x

OpenStudy (mertsj):

\[-2+\frac{3}{4}x=5x-\frac{1}{3}\]

OpenStudy (mertsj):

Multiply each term by 12: \[12\times(-2)+12\times\frac{3}{4}x=12\times(5x)-12(\frac{1}{3})\]

OpenStudy (mertsj):

\[-24+9x=60x-4\]

OpenStudy (mertsj):

You can finish it.

OpenStudy (anonymous):

yeah thank you